# GET: Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v1

@@ 4.2 Dual-branch multi-modal joint training
$\mathcal{L}_{\text{Dual}}=\mathcal{L}_{\text{dual}}^{v}+\mathcal{L}_{\text{dual}}^{t}-\lambda_{h}H_{mm}+\lambda_{c}\mathcal{L}_{\text{CICO}}\,.$ (15)
@@ 5.1 Experimental setup
Datasets. We evaluate our method on multiple benchmarks, including three image classification generic datasets (i.e., CIFAR 10/100 [20] and ImageNet-100 [8]), three fine-grained datasets from Semantic Shift Benchmark [37] (i.e., CUB [39], Stanford Cars [19] and FGVC-Aircraft [27]), and three challenging datasets (i.e., Herbarium 19 [35], ImageNet-R [15] and ImageNet-1K [8]). Notice that we are the first to introduce ImageNet-R into the GCD task, which contains various renditions of 200 ImageNet classes, thus challenging the GCD’s assumption that the data comes from the same domain. The details of the datasets we evaluate on are reported in Supplementary Materials (SM).
Evaluation metrics.
Following standard evaluation protocol (V2) in [36], we evaluate the performance with clustering accuracy (ACC), calculated as:
$\mathcal{ACC}=\max_{perm\in\mathcal{P}}\frac{1}{|\mathcal{D}_{u}|}\sum_{i=1}^{|\mathcal{D}_{u}|}\mathbbm{1}{[y_{i}=perm(\hat{y}_{i})]}\,,$ (16)
where $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted label, respectively. $\mathcal{P}$ is the set of all permutations while the optimal can be computed with the Hungarian algorithm [22].
Implementation details.
We use a ViT-B/16 [9] pre-trained with CLIP [32] as the backbone of the image and text encoder. In the first stage, we train a fully connected layer to transfer image embeddings into pseudo tokens. In the second stage, we use a single linear layer to turn pseudo text embeddings generated by TES into learnable embeddings while changing their dimensions to match those of the visual features. Our dual-branch training implementation details simply follow previous [40, 36] and are presented in SM. The last-epoch visual branch is used for evaluation and we provide the pseudocode in SM to clearly describe the details of the training and testing.
@@ 5.2 Comparison with state of the arts
In this section, we compare the proposed GET with the several state-of-the-art methods, including: $k$-means [26], RS+ [13], UNO+ [11], ORCA [2], PromptCAL [43], DCCL [31], GPC [46], GCD [36] and SimGCD [40]. GCD [36] and SimGCD [40] provide paradigms for non-parametric and parametric classification, thus we replace their backbone with CLIP for a fair comparison, denoted as GCD-CLIP and SimGCD-CLIP.
Evaluation on fine-grained datasets.
As shown in Tab. 1, our method achieves remarkable success on all three fine-grained datasets. Specifically, we surpass SimGCD-CLIP by 5.3%, 8.5%, and 4.6% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. In fine-grained datasets, the visual conceptions of distinct classes exhibit high similarity, making it challenging for classification based solely on visual information. However, their text information is notably different. Consequently, our GET  significantly enhances classification accuracy through the reciprocal augmentation of text and visual information flow. More qualitative results are reported in SM.
Table 1: Results on fine-grained and generic datasets.
@@
Evaluation on generic datasets.
In Tab. 1, we also present the performances for three generic datasets. Due to the low resolution of the CIFAR dataset and model biases(CLIP itself performs poorly on CIFAR100, with a zero-shot performance of 68.7), the results for novel classes are inferior compared to the DINO backbone. However, despite the inherent limitations in the discriminative capability of CLIP itself, our method still achieves an improvement of 0.4% on ‘Old’ classes of CIFAR10 and 2.2% on ‘New’ classes of CIFAR100, compared to SimGCD-CLIP. For ImageNet-100, SimGCD-CLIP has achieved an exceptionally saturated result of 90.8% on ‘All’ classes, further advancements pose considerable challenges. However, leveraging the additional modality information, GET elevates the performance ceiling to an impressive 91.7%.
Table 2: Results on more challenging datasets.
@@
Evaluation on challenging datasets.
We further evaluate on three challenging datasets. As shown in Tab. 2, GET outperforms all other state-of-the-art methods for both ‘All’ and ‘New’ classes on Herb19 and ImageNet-1K datasets. In particular, our method achieves a notable improvement of 1.4% and 1.7% on ‘New’ classes of Herb19 and ImageNet-1K, respectively. Furthermore, the suboptimal performance of GCD and SimGCD with the DINO backbone on the ImageNet-R dataset highlights the difficulty of DINO in discovering new categories with multiple domains. Despite multiple domains for images of the same category, their textual information remains consistent. Our method effectively integrates text information, resulting in a substantial improvement of 3.2% and 6.0% over the state-of-the-art for ‘All’ classes and ‘Old’ classes, respectively. It is worth noting that, owing to the text consistency within the same category, our text branch achieves remarkable 62.6% and 63.5% accuracy for ‘All’ classes of Imagenet-R and ImageNet-1K, respectively. The detailed results of the text branch can be seen in SM.
@@ 5.3 Ablation study and analysis
To evaluate the effectiveness of different components, we conduct an ablation study on Stanford Cars and CIFAR100 in Tab. 3. Comparing (2) with (1), the dual-branch training strategy can effectively leverage the text features generated by TES, resulting in a 6.2% improvement on Scars’ ‘All’ classes and a 0.3% improvement on CIFAR100’s ‘Old’ classes. Furthermore, comparing (3) with (1), CICO enables the two branches to exchange information and mutually benefit from each other, resulting in remarkable improvements of 11% on Scars’ ‘New’ and 2.2% on CIFAR100’s ‘New’.
Table 3: Ablation study of different components.
@@
Effectiveness of text embedding synthesizer.
In order to prove that our text embedding synthesizer can generate reliable and discriminative representations, we visualize the text embeddings of CIFAR10 with t-SNE. As shown in Fig. 3, the initial text embeddings within the same class exhibit clear clustering, and the learnable embeddings further produce compacter clusters. Moreover, we introduce TES into the non-parametric GCD by straightforwardly concatenating text and image features before semi-supervised k-means classification. As in Tab. 4, with the help of text information, GCD gains about 5% average improvement on ‘All’ classes over 3 datasets, demonstrating the importance of multi-modal information in GCD task and our TES can be widely used in multiple GCD methods. We provide more results and analysis for TES in SM.
Table 4: Effectiveness of TES in non-parametric GCD.
FGVC-Aircraft ImageNet-100 ImageNet-R
Methods All Old New All Old New All Old New
GCD-CLIP 45.3 44.4 45.8 75.8 87.3 70.0 44.3 79.0 25.8
+TES 49.6 49.3 49.8 80.0 95.1 72.4 49.4 79.4 33.5
Comparison with different fusion methods.
In Tab. 5, we compare our dual-branch training strategy with other methods that fuse visual and text features, including concatenation and mean. Although they may show improvements in ‘Old’ or ‘New’ classes due to the multi-modal information, we demonstrate that joint learning of the two branches is more effective as it encourages the model to have complementary and aligned discriminative capabilities for visual and text features of a same class, leading to more discriminative multi-modal prototypes and more cohesive multi-modal clusters.
Table 5: Comparison with different fusion methods.

### v2

@@ 4.2 Dual-branch multi-modal joint training
$\mathcal{L}_{\text{Dual}}=\mathcal{L}_{\text{db}}^{v}+\mathcal{L}_{\text{db}}^{t}-\epsilon H_{mm}+\lambda_{c}\mathcal{L}_{\text{CICO}}\,.$ (14)
Since information from different modalities is exchanged and learned through CICO and injected into the trainable visual backbone, we utilize the last-epoch visual branch for inference.
The training procedure of the proposed GET is presented in Algorithm 1.
Input: Training dataset $\mathcal{D}=\mathcal{D}_{l}\cup\mathcal{D}_{u}$, a FC layer $l(\cdot|\theta_{\text{t}})$ and a MLP layer $g(\cdot|\theta_{\text{m}})$, fixed CLIP’s image encoder $E_{i}$ and text encoder $E_{t}$, a trainable image encoder $f_{v}(\cdot|\theta_{\text{v}})$, a prototypical classifier $\eta(\cdot|\theta_{\text{c}})$ and a linear projection $p(\cdot|\theta_{\text{p}})$.
Output: Predicted label $\hat{y}_{i}$.
/* Stage 1: TES Training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
            $\boldsymbol{z}_{i}^{v}=E_{i}(\boldsymbol{x}_{i})$ // visual embedding
            $\boldsymbol{t}_{i}=l(\boldsymbol{z}_{i}^{v}))$ // pseudo text tokens
            $\boldsymbol{\hat{z}}_{i}^{t}=E_{t}(\boldsymbol{t}_{i})$ // pseudo text embedding
            $\mathcal{L}_{align}\leftarrow$Eqs. (5) and (6)
            $\mathcal{L}_{distill\leftarrow}$q. (7)
            $\mathcal{L}_{TES}\leftarrow$Eq. (8)
            Back-propagation and optimize $\theta_{\text{t}}$.
until reaching max epochs;
/* Stage 2: Dual-branch training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
             /* Visual-branch */
             $\boldsymbol{z}_{i}^{v}=f_{v}(\boldsymbol{x}_{i}),\ \boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v}),\ \boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$
            Compute $\mathcal{L}_{ucon}^{v}$ and $\mathcal{L}_{scon}^{v}$ by replacing $\boldsymbol{h}$ in Eqs.(1) and (2) with $\boldsymbol{h}^{v}$
            $\mathcal{L}_{\text{rep}}^{v}\leftarrow$Eq. (9)
            Compute $\mathcal{L}_{\text{cls}}^{v}$ by replacing $\boldsymbol{p}$ in Eqs.(3) and (4) with $\boldsymbol{p}^{v}$
            $\mathcal{L}_{\text{db}}^{v}\leftarrow$Eq. (10)
            /* text-branch */
             $\boldsymbol{\hat{z}}_{i}^{t}=E_{t}(l(E_{i}(\boldsymbol{x}_{i})))$
            $\boldsymbol{\hat{z}}_{i}^{tl}=p(\boldsymbol{\hat{z}}_{i}^{t}),\ \boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl}),\ \boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$
            $\mathcal{L}_{\text{db}}^{t}\leftarrow$Eq. (11)
            Compute the multi-modal mean entropy regularization $H_{mm}$
            /* CICO */
             Calculate the visual and text anchors $\mathcal{P}_{v},\mathcal{P}_{t}$
            Compute the instance relationships by Eq. (12)
            $\mathcal{L}_{\text{CICO}}\leftarrow$Eq. (13)
            $\mathcal{L}_{\text{Dual}}\leftarrow$Eq. (14)
            Back-propagation and optimize $\theta_{\text{v}},\ \theta_{\text{m}},\ \theta_{\text{c}},\ \theta_{\text{p}}$.
until reaching max epochs;
return $\hat{y}_{i}=\eta(f_{v}(\boldsymbol{x}_{i}))$.
Algorithm 1 Pseudocode for  GET.
@@ 5.1 Experimental setup
Datasets.  We evaluate our method on multiple benchmarks, including three image classification generic datasets (i.e., CIFAR 10/100 [31] and ImageNet-100 [32]), three fine-grained datasets from Semantic Shift Benchmark [33] (i.e., CUB [34], Stanford Cars [35] and FGVC-Aircraft [36]), and three challenging datasets (i.e., Herbarium 19 [37], ImageNet-R [38] and ImageNet-1K [32]). Notice that we are the first to introduce ImageNet-R into the GCD task, which contains various renditions of 200 ImageNet classes, thus challenging the GCD’s assumption that the data comes from the same domain. For ImageNet-R, we subsample the first 100 classes as old classes, leaving the rest as new classes; the labeled dataset $\mathcal{D}_{l}$ consists of half of the old class samples, while the other half and all the new class samples are used to construct unlabelled dataset $\mathcal{D}_{u}$. As for other benchmarks, we follow the previous [3, 4] to sample $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$. The details of the datasets we evaluate on are shown in Table I.
TABLE I: The details of the datasets we evaluate on.
Labelled Unlabelled
Dataset Images Classes Images Classes
CIFAR10 [31] 12.5K 5 37.5K 10
CIFAR100 [31] 20.0K 80 30.0K 100
ImageNet-100 [32] 31.9K 50 95.3K 100
CUB [34] 1.5K 100 4.5K 200
Stanford Cars [35] 2.0K 98 6.1K 196
FGVC-Aircraft [36] 1.7K 50 5.0K 100
Herbarium 19 [37] 8.9K 341 25.4K 683
ImageNet-R [38] 7.7K 100 22.3K 200
ImageNet-1K [32] 321K 500 960K 1000
Evaluation metrics.  Following standard evaluation protocol (V2) in [3], we evaluate the performance with clustering accuracy (ACC), calculated as:
$\mathcal{ACC}=\frac{1}{|\mathcal{D}_{u}|}\sum_{i=1}^{|\mathcal{D}_{u}|}\mathbbm{1}{[y_{i}=perm(\hat{y}_{i})]}\,,$ (15)
where $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted label, respectively. $perm$ is the optimal permutation to obtain the highest ACC and can be computed with the Hungarian algorithm [39].
Implementation details.  We use a ViT-B/16 [16] pre-trained with CLIP [9] as the backbone of the image and text encoder. In the first stage, we train a fully connected layer to transfer image embeddings into pseudo tokens. In the second stage, we only fine-tune the last transformer block of the image encoder for all datasets and remove the projector resulting in features with a dimension of 768. The exception is ImageNet-1K, we remain and fine-tune the last projection layer, which avoids gradient explosion and improves results with lower computational cost, resulting in features with a dimension of 512. We use a single linear layer to turn pseudo text embeddings generated by TES into learnable embeddings while changing their dimensions (512 to 768) to match those of the visual features. The batch size is fixed to 128 for training and 256 for testing. Training is done with an SGD optimizer and an initial learning rate of 0.1 decayed by a cosine annealing rule. We train for 200 epochs on each dataset in both two stages. In the first stage, we set the number of pseudo text tokens to 7. The balance coefficient $\lambda$ is set to 0.35 as [3], while $\lambda_{c}$ is set to 2 and 1, respectively. The temperature value $\tau_{a}$ is set to 0.01 while other temperature values $\tau_{c}$, $\tau_{s}$ and $\tau_{t}$ are as same as [4]. All experiments are conducted with 4 NVIDIA GeForce RTX 3090 GPUs.
@@ 5.2 Comparison with state of the arts
In this section, we compare the proposed GET with the several state-of-the-art methods, including: $k$-means [40], RS+ [12], UNO+ [14], ORCA [41], PromptCAL [5], DCCL [42], GPC [43], GCD [3] and SimGCD [4]. GCD [3] and SimGCD [4] provide paradigms for non-parametric and parametric classification, thus we reproduce them by replacing their backbone with CLIP for a fair comparison, denoted as GCD-CLIP and SimGCD-CLIP. The results presented in the paper are the averages of three independent repeated runs. We provide the performance standard deviation of main results in the Appendix.
Evaluation on fine-grained datasets.  As shown in Tab. II, our method achieves consistently remarkable success on all three fine-grained datasets. Specifically, we surpass SimGCD-CLIP by 5.3%, 8.5%, and 4.6% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. In fine-grained datasets, the visual conceptions of distinct classes exhibit high similarity, making it challenging for classification based solely on visual information. However, their text information can provide additional discriminative information. Consequently, our GET  significantly enhances classification accuracy through the reciprocal augmentation of text and visual information flow.
TABLE II: Results (%) on fine-grained and generic datasets. The best results are highlighted in bold
@@
TABLE III: Results (%) on more challenging datasets. The best results are highlighted in bold
@@
Evaluation on generic datasets.  In Table II, we also present the performances for three generic datasets. Due to the low resolution of the CIFAR dataset and model biases (CLIP itself performs poorly on CIFAR100, with a zero-shot performance of 68.7), the results for novel classes are inferior compared to the DINO backbone. However, despite the inherent limitations in the discriminative capability of CLIP itself, our method still achieves an improvement of 0.4% on ‘Old’ classes of CIFAR10 and 2.2% on ‘New’ classes of CIFAR100, compared to SimGCD-CLIP. For ImageNet-100, SimGCD-CLIP has achieved an exceptionally saturated result of 90.8% on ‘All’ classes, further advancements pose considerable challenges. However, leveraging the additional modality information, GET elevates the performance ceiling to an impressive 91.7%.
Evaluation on challenging datasets.  We further evaluate on three challenging datasets. As shown in Tab. III, GET outperforms all other state-of-the-art methods for both ‘All’ and ‘New’ classes on Herb19 and ImageNet-1K datasets. In particular, our method achieves a notable improvement of 1.4% and 1.7% on ‘New’ classes of Herb19 and ImageNet-1K, respectively. Furthermore, the suboptimal performance of GCD and SimGCD with the DINO backbone on the ImageNet-R dataset highlights the difficulty of DINO in discovering new categories with multiple domains. Despite multiple domains for images of the same category, their textual information remains consistent. Our method effectively integrates text information, resulting in a substantial improvement of 3.2% and 6.0% over the state-of-the-art for ‘All’ classes and ‘Old’ classes, respectively. It is worth noting that, owing to the text consistency within the same category, our text branch achieves remarkable 62.6% and 63.5% accuracy for ‘All’ classes of Imagenet-R and ImageNet-1K, respectively. The detailed results of the text branch can be seen in the Appendix.
@@ 5.3 Ablation study and analysis
TABLE IV: Ablation study of different components.
@@
To evaluate the effectiveness of different components, we conduct an ablation study on Stanford Cars and CIFAR100 in Tab. IV. Comparing (2) with (1), the dual-branch training strategy can effectively leverage the text features generated by TES, resulting in a 6.2% improvement on Scars’ ‘All’ classes and a 0.3% improvement on CIFAR100’s ‘Old’ classes. Furthermore, comparing (3) with (1), CICO enables the two branches to exchange information and mutually benefit from each other, resulting in remarkable improvements of 11% on Scars’ ‘New’ and 2.2% on CIFAR100’s ‘New’.
Comparison with different fusion methods.  In Tab. V, we compare our dual-branch training strategy with other methods that fuse visual and text features, including concatenation and mean. Although they may show improvements in ‘Old’ or ‘New’ classes due to the multi-modal information, we demonstrate that joint learning of the two branches is more effective as it encourages the model to have complementary and aligned discriminative capabilities for visual and text features of a same class, leading to more discriminative multi-modal prototypes and more cohesive multi-modal clusters.
TABLE V: Comparison with different fusion methods.